# TopoNav: Topological Navigation for Efficient Exploration in Sparse Reward Environments

### v2

@@ III Background and Problem Formulation
In this section, we provide our problem formulation and briefly explain the topological mapping used for navigation and the Hierarchical Deep Q-Networks (h-DQN) used as a basis for our policy learning.
III-A Problem Formulation
@@
where $\pi^{*}$ is the optimal policy, $\pi_{h}:\mathcal{S}\rightarrow\mathcal{A}_{h}$ is the high-level policy (meta-controller), $\pi_{l}:\mathcal{S}\times\mathcal{A}_{h}\rightarrow\mathcal{A}_{l}$ is the low-level policy (controller), $r_{t}$ is the reward at time step $t$, and $\gamma$ is the discount factor.
III-B Topological Mapping for Navigation
Topological mapping represents the environment as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $v\in\mathcal{V}$ correspond to distinct places or landmarks, and edges $e\in\mathcal{E}$ represent the connectivity between them. This representation allows for efficient path planning and navigation by abstracting away the metric details and focusing on the high-level structure of the environment  [15]. In our approach, we dynamically construct and update the topological map $\mathcal{M}$ as the robot explores the environment. Each node $v_{i}\in\mathcal{V}$ is associated with a feature vector $\mathbf{f}i$ that encodes the sensory observations at that location. The edges $e_{ij}\in\mathcal{E}$ represent the traversability between nodes $v_{i}$ and $v_{j}$, which can be determined based on the robot’s motion model and the observed environmental conditions. The topological map $\mathcal{M}$ is used to guide the exploration process and enable efficient decision-making. By representing the environment as a graph, the robot can plan paths between different nodes and make high-level decisions based on the connectivity of the map. This allows for more strategic exploration and navigation compared to purely reactive approaches.
III-C Hierarchical Deep Q-Networks (H-DQN)
@@ IV-A Attention-based Feature Detection
TopoNav integrates an attention-based feature detection module, combining a ResNet-50 CNN [29] with a Convolutional Block Attention Module (CBAM) [30], to detect landmarks, objects, and trees from RGB images $\mathbf{I}_{i}\in\mathbb{R}^{H\times W\times 3}$. This module processes images to produce feature maps $\mathbf{F}_{i}\in\mathbb{R}^{h\times w\times c}$, refining them through CBAM to emphasize relevant features for landmark detection. Utilizing a sliding window approach on the refined feature map $\mathbf{F}^{\prime}_{i}$, TopoNav identifies potential Regions of Interest (ROIs) corresponding to navigational markers. Each ROI’s feature vector $\mathbf{f}_{ROI}\in\mathbb{R}^{d}$, obtained through ROI pooling, undergoes a binary classification to ascertain the presence of valid landmarks, objects, or trees, assigning a probability score $p_{ROI}\in[0,1]$. For each detected landmark, object, or tree, TopoNav extracts its corresponding feature vector $\mathbf{f}_{ROI}$ and uses it to create a new node or subgoal in the topological map.
@@ 11             $l_{\text{best}}=l$
During navigation, the meta-controller selects the next subgoal based on the current state and the robot’s overall goal. If no landmarks are detected within a certain distance threshold, TopoNav generates a new subgoal along the robot’s current trajectory at the maximum detection range to ensure continuous progress towards the final goal. The sub-controller’s policy generates a sequence of primitive actions or low-level skills to navigate the robot towards the selected subgoal, aiming to maximize the expected cumulative reward. The navigation process continues until the robot reaches the final goal or a maximum number of steps is exceeded. (See Algorithm  2)
@@ IV-D Hierarchical Policy Learning
Theorem 1 (Subgoal Reachability). Given a topological map $G=(V,E)$ constructed by TopoNav, with $V$ representing nodes (subgoals) and $E$ representing edges (navigable paths), for any two nodes $v_{i},v_{j}\in V$, if a path exists from $v_{i}$ to $v_{j}$ in $G$, then the sub-controller policy $\pi_{s}$ learned by TopoNav can navigate the robot from $v_{i}$ to $v_{j}$ with a probability of at least $1-\epsilon$. Here, $\epsilon$ is a small positive constant representing the upper bound of navigation failure probability, which accounts for environmental stochasticity and the convergence of the learning algorithm.
@@ V-A Evaluation Environments
The TopoNav framework is evaluated in a range of simulated and real-world environments. The simulated environments are created using the Unity 3D engine and include outdoor urban scenes, and off-road terrains. The real-world experiments are conducted in an outdoor off-road environment using a Clearpath Jackal robot. The environments are designed to cover various navigation challenges, such as narrow passages, obstacles, uneven terrain, and dead ends. The size of the environments ranges from 20m × 20m to 200m × 200m, and the complexity is varied by adjusting the density and arrangement of obstacles. For each environment, multiple starting and goal locations are randomly sampled to create a diverse set of navigation tasks. The performance of TopoNav is evaluated over 100 episodes in each environment, and the average metrics are reported.
@@ V-C Baselines
Figure 3: Fig. 3: (a-d) The robot navigates through various outdoor environments: (a) an open space without obstacles/vegetation (Scenario 1), (b) a natural setting with trees/vegetation (Scenario 2), (c) a cluttered environment with obstacles and landmarks (Scenario 3), and (d) a diverse terrain in simulation. (e-h) The corresponding topological maps generated by TopoNav for each scenario. The topological maps capture the connectivity and traversability of the environments, representing key locations and paths. The green dots in the maps represent the nodes or subgoals, which correspond to landmarks or distinct places. The edges (red lines) indicate the navigability between these nodes, enabling efficient path planning and navigation for the robot in each setting.
@@ V-D Implementation Details
Our RL network is implemented in PyTorch and trained using simulated terrains with a Clearpath Husky robot in ROS Noetic and Unity Simulation framework, which allows for testing in realistic outdoor environments with diverse terrain and objects. The simulated Husky robot is mounted with a Velodyne VLP16 3D LiDAR. The network is trained in a workstation with a 10th-generation Intel Core i9-10850K processor and an NVIDIA GeForce RTX 3090 GPU. For real-time deployment and inference, we use the Jackal UGV from Clearpath Robotics equipped with a 3D VLP-32C Velodyne Ultrapuck LiDAR and an AXIS Fixed IP Camera. It also includes an onboard Intel computer system, equipped with an Intel i7-9700TE CPU and an NVIDIA GeForce GTX 1650 Ti GPU. We utilize the LiDAR sensor provided 3D point cloud data for obstacle detection and avoidance using the Dynamic Window Approach (DWA)  [32].
@@ Acknowledgment
This work has been partially supported by U.S. Army Grant #W911NF2120076

### v3

@@ III Background and Problem Formulation
In this section, we outline the topological mapping for navigation, the Hierarchical Deep Q-Networks (H-DQN) for policy learning, and our problem formulation.
III-A Topological Mapping
Topological mapping represents the environment as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $v\in\mathcal{V}$ correspond to distinct places or landmarks, and edges $e\in\mathcal{E}$ represent the connectivity between them. This representation allows for efficient path planning and navigation by abstracting away the metric details and focusing on the high-level structure of the environment  [15]. In our approach, we dynamically construct and update the topological map $\mathcal{M}$ as the robot explores the environment. Each node $v_{i}\in\mathcal{V}$ is associated with a feature vector $\mathbf{f}_{i}$ that encodes the sensory observations at that location. The edges $e_{ij}\in\mathcal{E}$ represent the traversability between nodes $v_{i}$ and $v_{j}$, which can be determined based on the robot’s motion model and the observed environmental conditions. The topological map $\mathcal{M}$ is used to guide the exploration process and enable efficient decision-making. By representing the environment as a graph, the robot can plan paths between different nodes and make high-level decisions based on the connectivity of the map. This allows for more strategic exploration and navigation compared to purely reactive approaches.
III-B Hierarchical Deep Q-Networks (H-DQN)
III-C Problem Formulation
@@
where $\pi^{*}$ is the optimal policy, $\pi_{h}:\mathcal{S}\rightarrow\mathcal{A}_{h}$ is the high-level policy (meta-controller), $\pi_{l}:\mathcal{S}\times\mathcal{A}_{h}\rightarrow\mathcal{A}_{l}$ is the low-level policy (controller), $r_{t}$ is the reward at time step $t$, and $\gamma$ is the discount factor. In this context, the state space $\mathcal{S}$ includes a dynamically constructed topological map $\mathcal{M}$, where nodes represent automatically identified landmarks. The high-level action space $\mathcal{A}_{h}$ consists of subgoals, allowing the meta-controller to make decisions based on the evolving map. This approach enables efficient navigation in unknown environments without relying on predefined waypoints.
@@ IV-A Attention-based Feature Detection
TopoNav integrates an attention-based feature detection module, combining a ResNet-50 CNN [29] with a Convolutional Block Attention Module (CBAM) [30], to detect landmarks, objects, and trees from RGB images $\mathbf{I}_{i}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ are the height and width of the input image, respectively. This module processes images to produce feature maps $\mathbf{F}_{i}\in\mathbb{R}^{h\times w\times c}$ ($h$ and $w$ are the height and width of the feature map, and $c$ is the number of channels) refining them through CBAM to emphasize relevant features for landmark detection. Utilizing a sliding window approach on the refined feature map $\mathbf{F}^{\prime}_{i}$, TopoNav identifies potential Regions of Interest (ROIs) corresponding to navigational markers. Each ROI’s feature vector $\mathbf{f}_{ROI}\in\mathbb{R}^{d}$, obtained through ROI pooling, undergoes a binary classification to ascertain the presence of valid landmarks, objects, or trees, assigning a probability score $\mathbf{P}_{ROI}\in[0,1]$. For each detected landmark, object, or tree, TopoNav extracts its corresponding feature vector $\mathbf{f}_{ROI}$ and uses it to create a new node or subgoal in the topological map.
@@ 11             $l_{\text{best}}=l$
[…]
Figure 3: An illustration of the topological navigation process using detected landmarks as sub-goals and strategic landmark selection.
@@
During navigation, the meta-controller selects the next subgoal based on the current state and the robot’s overall goal. If no landmarks are detected within a certain distance threshold, TopoNav generates a new subgoal along the robot’s current trajectory at the maximum detection range to ensure continuous progress towards the final goal. The sub-controller’s policy generates a sequence of primitive actions or low-level skills to navigate the robot towards the selected subgoal, aiming to maximize the expected cumulative reward. The navigation process continues until the robot reaches the final goal or a maximum number of steps is exceeded. (See Algorithm  2 & Fig. 3)
@@ IV-D Hierarchical Policy Learning
Theorem 1 (Subgoal Reachability). Given a topological map $G=(V,E)$ constructed by TopoNav, with $V$ representing nodes (subgoals) and $E$ representing edges (navigable paths), for any two nodes $v_{i},v_{j}\in V$, if a path exists from $v_{i}$ to $v_{j}$ in $G$, then the sub-controller policy $\pi_{s}$ learned by TopoNav can navigate the robot from $v_{i}$ to $v_{j}$ with a probability of at least $1-\epsilon$. Here, $\epsilon$ is a small positive constant representing the upper bound of navigation failure probability, which accounts for environmental stochasticity and the convergence of the learning algorithm. TopoNav minimizes $\epsilon$ through robust feature extraction and continuous map refinement.
@@ V-A Evaluation Environments
The TopoNav framework is evaluated in a range of simulated and real-world environments. The simulated environments are created using the Unity 3D engine and include outdoor urban scenes and off-road terrains. The simulated environments incorporate landmarks and objects similar to those in our real-world test scenarios (Fig. 4(d)). The real-world experiments are conducted in an outdoor off-road environment using a Clearpath Jackal robot. The environments are designed to cover various navigation challenges, such as narrow passages, obstacles, uneven terrain, and dead ends. The size of the environments ranges from 20m × 20m to 200m × 200m, and the complexity is varied by adjusting the density and arrangement of obstacles. For each environment, multiple starting and goal locations are randomly sampled to create a diverse set of navigation tasks. The performance of TopoNav is evaluated over 100 episodes in each environment, and the average metrics are reported.
@@ V-C Baselines
Figure 4: (a-d) The robot navigates through various outdoor environments: (a) an open space without obstacles/vegetation (Scenario 1), (b) a natural setting with trees/vegetation (Scenario 2), (c) a cluttered environment with obstacles and landmarks (Scenario 3), and (d) a diverse terrain in simulation. (e-h) The corresponding topological maps generated by TopoNav for each scenario. The topological maps capture the connectivity and traversability of the environments, representing key locations and paths. The green dots in the maps represent the nodes or subgoals, which correspond to landmarks or distinct places. The edges (red lines) indicate the navigability between these nodes, enabling efficient path planning and navigation for the robot in each setting.
@@ V-D Implementation Details
Our RL network is implemented in PyTorch and trained using simulated terrains with a Clearpath Husky robot in ROS Noetic and Unity Simulation framework, which allows for testing in realistic outdoor environments with diverse terrain and objects. The simulated Husky robot is mounted with a Velodyne VLP16 3D LiDAR. The network is trained in a workstation with a 10th-generation Intel Core i9-10850K processor and an NVIDIA GeForce RTX 3090 GPU. During training, our network uses a batch size of 128 and performs gradient updates using the Adam optimizer with learning rate $\lambda=10^{-4}$. The training process occurs over 1 million steps in our simulation environment, with performance evaluated at regular intervals. For real-time deployment and inference, we use the Jackal UGV from Clearpath Robotics equipped with a 3D VLP-32C Velodyne Ultrapuck LiDAR and an AXIS Fixed IP Camera. It also includes an onboard Intel computer system, equipped with an Intel i7-9700TE CPU and an NVIDIA GeForce GTX 1650 Ti GPU. We utilize the LiDAR sensor provided 3D point cloud data for obstacle detection and avoidance using the Dynamic Window Approach (DWA)  [32].
@@ Acknowledgment
This work has been partially supported by the DEVCOM Army Research Laboratory (ARL) under a cooperative agreement (W911NF2120076), ONR Grant #N00014-23-1-2119, NSF CAREER Award #1750936, NSF REU Site Grant #2050999, and NSF CNS EAGER Grant #2233879.